# Zero- to low-field -spectroscopy with a diamond magnetometer

### v1

@@ III.4 Experimental arrangement
The setup is sketched in Fig. 2. The experiment is conducted inside an MS-2 magnetic shield from Twinleaf. The OPM sensor is mounted within a 3D-printed holder; it is sensitive along the same axis as the diamond sensor (x-axis). OPM and diamond sensor are positioned on the x-axis around the spherical NMR sample (12.5 mm diameter). Both sensors are inside a Helmholtz-coil pair oriented along the sensing direction which provides the excitation pulse to initiate the free decay of the NMR sample magnetization. The NMR sample is enclosed within a solenoid wound around a ceramic tube (a “piercing” solenoid) piercingsolenoid that allows applying magnetic fields along the y-axis without disturbing the OPM sensor. This arrangement allows observation of nuclear spin precession around the y-axis within the x-z plane during the field precession experiments.
@@ V Discussion & Conclusion
The high bandwidth of the diamond sensor is useful to study -spectra of molecules with large -coupling. Platinum -couplings, for example, have been calculated to lie in the range of tens of kHz platinumnmr and PF exhibits - couplings of several multiples of 700 Hz pf6; Fabricant2024.
Similarly, the study of the NMR spectra in the intermediate or uncoupled regimes is easier accessible with higher bandwidth, providing the most complete information about the spin system (including both -couplings and chemical shifts) phip. We demonstrate the advantage of the higher bandwidth of the diamond sensor in Fig. 4 f). At fields where the precession frequency exceeds 500 Hz no proper response from the OPM can be obtained as it goes beyond the 6th order hardware low-pass filter quspin.
In the following we discuss the generalization of our results to other ZULF NMR samples. Firstly, at zero field a source of polarization is always required to produce detectable signals BARSKIY2025101558. Among the various techniques available Eills2023SpinHyperpolarization, Signal Amplification by Reversible Exchange (SABRE) is selected in this study due to its ability to produce high molar polarization under ambient conditions. However, the methodology presented here is not limited to SABRE. One can also employ several complementary hyperpolarization methods, including high-field prepolarization followed by sample shuttling, dynamic nuclear polarization (DNP), photo-induced DNP, and other parahydrogen-based approaches. While SABRE enables the first demonstration to our knowledge of diamond-based ZULF NMR, these alternative strategies broaden the scope of detectable molecular targets, including those not amenable to parahydrogen-based hyperpolarization. Thus, we view the use of NV sensors in conjunction with diverse hyperpolarization techniques as a viable platform for scalable and versatile low-field NMR detection for a variety of polarizable compounds.
@@ VII.1 Diamond sensor design choices
Pictures of the sensor head and the arrangement inside the magnetic shield can be seen in Fig. 7 and the most important components of the sensor head are detailed in Fig. 8. The diamond optics used for the collection of light are described in detail in reference diamondoptics. The microwave (MW) and photodiode printed circuit board (PCB) designs are depicted in Fig. 6. The flexible PCB for the MW delivery has a thickness of m and a center hole for the sensing diamond to stick out by approximately m. The microwave board design was chosen such that circularly polarized microwaves could be applied. The photodiode PCB was designed to maximize compactness and avoid components that could generate residual magnetic fields. The photodiodes are mostly nonmagnetic (Hamamatsu S13228-01). Titanium screws (TI AL GmbH) were used to fasten the aluminum components. The fiber holder and the optics-holder rod were manufactured from titanium and aluminum, respectively.

### v2

@@ III.4 Experimental arrangement
The setup is sketched in Fig. 2. The experiment is conducted inside an MS-2 magnetic shield from Twinleaf. The OPM sensor is mounted within a 3D-printed holder; it is sensitive along the same axis as the diamond sensor (x-axis). OPM and diamond sensor are positioned on the x-axis around the spherical NMR sample (12.5 mm diameter). Both sensors are inside a Helmholtz-coil pair oriented along the sensing direction which provides the excitation pulse to initiate the free decay of the NMR sample magnetization. The NMR sample is enclosed within a solenoid wound around a ceramic tube (a “piercing” solenoid) Yashchuk et al. (2004) that allows applying magnetic fields along the y-axis without disturbing the OPM sensor. This arrangement allows observation of nuclear spin precession around the y-axis within the x-z plane during the field precession experiments.
@@ V Discussion & Conclusion
The high bandwidth of the diamond sensor is useful to study -spectra of molecules with large -coupling. Platinum -couplings, for example, have been calculated to lie in the range of tens of kHz Ducati,Lucas C. Marchenko and Autschbach,Jochen (2016) and PF exhibits - couplings of several multiples of 700 Hz Blanchard et al. (2021a); Fabricant et al. (2024).
Similarly, the study of the NMR spectra in the intermediate or uncoupled regimes is easier accessible with higher bandwidth, providing the most complete information about the spin system (including both -couplings and chemical shifts) Colell,Johannes et al. (2013). We demonstrate the advantage of the higher bandwidth of the diamond sensor in Fig. 4 f). At fields where the precession frequency exceeds 500 Hz no proper response from the OPM can be obtained as it goes beyond the 6th order hardware low-pass filter M. Zhao and Fujimoto .
In the following we discuss the generalization of our results to other ZULF NMR samples. Firstly, at zero field a source of polarization is always required to produce detectable signals Barskiy et al. (2025). Among the various techniques available Eills et al. (2023), Signal Amplification by Reversible Exchange (SABRE) is selected in this study due to its ability to produce high molar polarization under ambient conditions. However, the methodology presented here is not limited to SABRE. One can also employ several complementary hyperpolarization methods, including high-field prepolarization followed by sample shuttling, dynamic nuclear polarization (DNP), photo-induced DNP, and other parahydrogen-based approaches. While SABRE enables the first demonstration to our knowledge of diamond-based ZULF NMR, these alternative strategies broaden the scope of detectable molecular targets, including those not amenable to parahydrogen-based hyperpolarization. Thus, we view the use of NV sensors in conjunction with diverse hyperpolarization techniques as a viable platform for scalable and versatile low-field NMR detection for a variety of polarizable compounds.
@@ VII.1 Diamond sensor design choices
Pictures of the sensor head and the arrangement inside the magnetic shield can be seen in Fig. 7 and the most important components of the sensor head are detailed in Fig. 8. The diamond optics used for the collection of light are described in detail in reference Omar,Muhib et al. (2023). The microwave (MW) and photodiode printed circuit board (PCB) designs are depicted in Fig. 6. The flexible PCB for the MW delivery has a thickness of m and a center hole for the sensing diamond to stick out by approximately m. The microwave board design was chosen such that circularly polarized microwaves could be applied. The photodiode PCB was designed to maximize compactness and avoid components that could generate residual magnetic fields. The photodiodes are mostly nonmagnetic (Hamamatsu S13228-01). Titanium screws (TI AL GmbH) were used to fasten the aluminum components. The fiber holder and the optics-holder rod were manufactured from titanium and aluminum, respectively.